OpenStudy (anonymous):

Solve the equation for y. 5x+y=7

jimthompson5910 (jim_thompson5910):

Subtract 5x from both sides

OpenStudy (anonymous):

wen i subtract 5x from both sides i get y=2x

jimthompson5910 (jim_thompson5910):

you don't combine 7 and -5x to get 2x

jimthompson5910 (jim_thompson5910):

you just leave it as 7 - 5x or -5x + 7

OpenStudy (anonymous):

5(0) + y = 7 y=7 (0,7)

jimthompson5910 (jim_thompson5910):

So the final answer is y = 7 - 5x or y = -5x + 7